OpenStudy (anonymous):

how many sixths are in 3 1/3

OpenStudy (mertsj):

\[3 \frac{1}{3}=\frac{10}{3}=\frac{ ?}{6}\]

OpenStudy (mertsj):

Can you fill in the blank?

hero (hero):